OpenStudy (anonymous):

14. Simplify 3! (1 point) 2 5 3 6 15. Simplify 8P3. (1 point) 42 336 40,432 56 16. Simplify 15C3. (1 point) 182 455 2,730 910

OpenStudy (anonymous):

Hey :) Welcome to OpenStudy :)

OpenStudy (anonymous):

14. Simplify 3! (1 point) \[3 \times 2 \times 1 \] --can you do this?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

15. Simplify 8P3. (1 point) --this is a permutation \[P(n,r) = nPr = \frac{ n! }{ (n-r)! }\] --so 8 is n & the 3 is the r --plug them into the formula: \[\frac{ 8! }{ (8-3)! }\]

OpenStudy (anonymous):

yes #14.) is 6

OpenStudy (anonymous):

thanks for 14

OpenStudy (anonymous):

16. Simplify 15C3. (1 point) -- this is a combination \[C(n,r) = nCr = \frac{ n! }{ r! (n -r)! }\] -- so the 15 is the n & the 3 is the r --plug them into the formula: \[\frac{ n! }{ 3! (15 -3)! }\]

OpenStudy (anonymous):

i dont get 15 and 16

OpenStudy (anonymous):

oops * forgot the n \[\frac{ 15! }{ 3! (15 -3)! }\]

OpenStudy (anonymous):

15 would be 336 right?

OpenStudy (anonymous):

15. \[ \frac{ 8! }{ (8-3)! }\] \[ \frac{ 8! }{ 5! }\] \[ \frac{ 8\times7\times6\times5\times4\times3\times2\times1 }{ 5\times4\times3\times2\times1 }\] --cancel out from the 5 and after. --so ur left with \[8\times7\times6\] --what does this equal?

OpenStudy (anonymous):

so 15.) is 336 your right :)

OpenStudy (anonymous):

wb 16

OpenStudy (anonymous):

have you tried it?

OpenStudy (anonymous):

yes but its complicated

OpenStudy (anonymous):

\[ \frac{ 15! }{ 3! (15 -3)! }\] \[ \frac{ 15! }{ 3! (12)! }\] \[ \frac{ 15\times14\times13\times \times12! }{ 3\times2\times1 \times 12! }\] --the 12! cancel so what are you left with?

OpenStudy (anonymous):

--sorry for the double x x --its only one x

OpenStudy (anonymous):

answer is 455

OpenStudy (anonymous):

yes it is 455 :)

OpenStudy (anonymous):

help with this please